If the number of turns per unit length of a coil of solenoid is doubled, the self-inductance of the solenoid will


(a)    remain unchanged


(b)    be halved


(c)    be doubled


(d)    become four times


Grade:11

2 Answers

Hrishant Goswami
10 years ago

(d)

        Self inductance of a solenoid =(µAn^2)/l

        So, self induction  proportional to n2

So, inductance becomes 4 times when n is doubled.

vedha prabhu
5 years ago
MAKE ALL THE PARAMETERS CONSTANT AS 1ASSUME THAT 
L IS DIRECTLY PROPORTIONAL TO N^2 i.e L =N^2
IF N=4 THEN L=4^2 = 16(x)
IF2N=2x4=8 THEN L= 8^2 =64(4x)
IF N IS DOUBLED THEN L WILL BE QUAD